Circular No. 7227

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1999db, 1999dc, 1999dd, 1999de A. Gal-Yam and D. Maoz, Tel Aviv University, report for the Wise Ob-

servatory Optical Transients Search (cf. IAUC 6917, 7055) their discovery of four apparent supernovae on unfiltered CCD images taken by Gal-Yam (with A. Daniel on July 11) using the Wise Observatory 1-m telescope:

SN  1999 UT  α_{2000}  δ_{2000}  R  Offset
1999db July 8 23^{h}33^{m}03.^{s}64 + 5^{o}32′23′′.5 18.7  —
1999dc July 10 14 44 08.80 +58 53 31.6 19.3  —
1999dd July 10 0 27 31.06 +21 42 47.0 21.1  —
1999de July 11 0 22 28.20 +23 23 56.2 21.2 4′′.0 W, 4′′.8 N

SNE 1999db, 1999dc, and 1999de are located in the fields of the galaxy clusters Abell 2616 (z = 0.183), Abell 1966 (z = 0.151), and Abell 24 (z = 0.134), respectively. SNe 1999db, 1999dc, and 1999dd are superimposed on compact host galaxies, so no offsets can be determined from the images. SN 1999db was confirmed on July 9 by Gal-Yam, but was invisible in images obtained on 1998 Dec. 20. SN 1999dc was confirmed by Gal-Yam and Daniel on July 12, but was invisible in images obtained on June 13. SN 1999dd was confirmed by J. Dan and P. Ibbetson on July 14, but was invisible in images obtained on 1998 July 20. SN 1999de was confirmed by Dan and Ibbetson on July 14, but was invisible in images obtained on 1998 July 23. The limiting magnitude of all images is R = 22.0. Finding charts and further details can be obtained via http://wise-

obs.tau.ac.il/~avishay/snlist.html.

GX 17+2 S. F. Anderson, B. Margon, and E. W. Deutsch, University of Wash-

ington, write: “Archival K-band images of the GX 17+2 field obtained on 1998 Aug. 27 with the ARC 3.5-m telescope show an object blended (in modest 1′′ seeing) with NP Ser, at a position 1′′ north of, and only slightly fainter than, NP Ser itself — in good agreement with the recent report of the June 1999 bright state (IAUC 7219). This adds yet further weight to the identification of the x-ray source with star A originally suggested by Deutsch et al. (Ap.J., in press) from 1997 observations when the object was very faint and indicates that the bright state of the source first reported on IAUC 7219 might not necessarily be a brief or rare event.”

1999 July 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green